# Turning the Tide on Dark Pools? Towards Multi-Stakeholder Vulnerability Notifications in the Ad-Tech Supply Chain

### v1

@@ 6.2 Limitations
Multi-stakeholder interdependence. The online advertising multi-stakeholder supply chain is fundamentally interdependent with mixed incentives for each participant. Publishers and advertisers cannot directly fix their participation in dark pooling without cooperation from ad-networks. Ad-networks can address their own facilitation of dark pools but cannot control the behavior of other ad-networks using their platform. Using a single round notification (notifying all stakeholders simultaneously) would make it impossible to identify which entity resolved the vulnerability. To solve this, we used a multi-round notification design, notifying one stakeholder in each round. This approach allows us to identify more confidently which stakeholder resolved the vulnerability. However, there is still a risk of contamination when a stakeholder responds much later, after the next round has begun. We minimized such risks by leaving 4-5 weeks between each round of notifications, giving each stakeholder ample time to respond and remediate. Based on the email responses, we observed that nearly all arrived within two weeks of the initial notification. Additionally, publishers may reach out to ad-networks to remediate the vulnerability, before we notify ad-networks in the second phase, alerting them. However, publishers are incentivized to work with more ad-networks and vice versa to maximize their respective revenue. Hence, we acknowledge that complete isolation is not possible. However, since we detect vulnerabilities before each phase, ad-networks were notified of different or unresolved entities than the ones in notifications to publishers, guaranteeing confidence in the results we observe. We believe these approaches effectively addressed the challenge of interdependent multi-stakeholder notification research.
@@
Source of activist notifications. We sent our activist notifications in collaboration with the CheckMyAds Institute, a well-known organization in the ad-tech community. Due to operational constraints, we could only use their branding on our emails, reports, and the website. We could not use their email servers to send notifications due to risks to their server reputation and potential interference with their operations. This might have influenced how recipients perceived the email. To minimize this impact, we clearly stated CheckMyAds’ involvement in the opening sentence of the notification and used their branding prominently in the headers and footers of our emails, reports, and website.
6.3 Ethical considerations
@@
IRB review. Our institutional review board (IRB) reviewed this notification study and deemed it as not involving human subjects. Despite this, we designed and conducted our measurements and notifications following the principle of beneficence outlined in the Menlo Report [70] and Belmont Report [71]. We aimed to maximize benefits and minimize potential harms.
@@
Infrastructure costs and risks. We measured ad-tech supply vulnerabilities through crawls. To avoid stressing web servers, we did not conduct dynamic crawls of problematic publishers concurrently. We spaced out our additional crawls for ads.txt and sellers.json files between 12 days to one month apart. Our crawlers did not follow the robots.txt directives on problematic publishers, but our methodology aligns with ethical and legal considerations for such audits [72]. We sent notifications from a dedicated email server, set up with our university’s IT staff, to protect the reputation of the university’s mail servers.
Privacy risks. We did not collect or record any personal information in this research. Although we had phone calls and virtual meetings with several organizations to help them remediate vulnerabilities, we did not record or share any conversations or emails with anyone other than the authors.
@@
Advertising costs. To understand which brands advertise on problematic publishers and which ad-networks are responsible, we clicked on the ads shown during page loads. The costs associated with our 28,376 ad clicks were relatively minor, as per-ad CPMs are low, particularly for browsers without profiles [73, 74]. These costs are justifiable given the benefits of understanding and remedying supply-chain vulnerabilities in the ad-tech ecosystem.
@@
6.4 Concluding remarks
Implications for online advertising. Our research illuminates how the online advertising supply chain functions and how different stakeholders interact. Our qualitative analysis shows that notifications can raise awareness about vulnerabilities in the supply chain, leading to increased transparency, trust, and the termination of relationships with problematic actors. While notifications can effectively address vulnerabilities like dark pools in the online advertising ecosystem, there are important nuances. The motivations and capabilities of notification recipients play a crucial role. We found that publishers have the lowest dark pool remedial rates among stakeholders. In contrast, ad-networks, with their capability and resources, and advertisers, with their monetary incentives, are significantly more effective at resolving dark pools. Additionally, we found that the source of the notification, whether from activists or academics, does not affect dark pool resolution, except for publishers. Future work can investigate other notification channels (e.g., public naming and shaming [75]) and external stakeholders (e.g., direct consumer notifications [76]).
Implications for other supply chains. Our results highlight the need to consider the complexities of multiple stakeholders in other ecosystems, such as the software supply chain, which includes operators, administrators, vendors, and developers. This need is further highlighted by Zimmermann et al. [77] who showed that vulnerabilities in widely-used open-source libraries impacted numerous software products. Unfortunately, the current state-of-the-art in security and privacy vulnerability notification only considers singlular stakeholders. Similarly, in fintech, Abdou et al. [78] showed that inadequate API security could expose critical consumer financial data, impacting multiple stakeholders including consumers, banks, developers, and regulators in open banking ecosystem. Our work provides a model for testing the effectiveness of vulnerability notifications in these multi-stakeholder ecosystems.
Multi-stakeholder notification research warrants more attention. Research on multi-stakeholder supply chains is complicated due to stakeholder inter-dependencies and mixed incentives. These complexities raise questions about which stakeholder to notify and how to notify each one. Our study, using the online advertising supply chain as a case-study, shows that not all stakeholders are equally capable or willing to resolve vulnerabilities. Furthermore, changing notification sources only, significantly influences some stakeholders. This research is the first to tackle this scenario and highlights the need to examine other multi-stakeholder ecosystems.
Appendix
(a) Approaches to resolution
(b) Attitude towards notification
(c) Challenges to resolve vulnerability
(d) Approaches to resolution
(e) Attitude towards notification by
(f) Challenges to resolve vulnerability
(g) Approaches to resolution
(h) Attitude towards notification
(i) Challenges to resolve vulnerability
@@ 6.4 Concluding remarks
Figure 5: Categorization of notification responses by sender reputation under different themes for victim publishers [(a)–(c)], ad-networks [(d)–(e)], and advertisers [(f)–(h)].

### v2

@@ 6.2 Limitations
Multi-stakeholder interdependence. The online advertising multi-stakeholder supply chain is fundamentally fragmented – interdependent with mixed incentives for each participant. Publishers and advertisers cannot directly fix their participation in dark pooling without cooperation from ad-networks. Ad-networks can address their own facilitation of dark pools but cannot control the behavior of other ad-networks using their platform. Using a single round notification (notifying all stakeholders simultaneously) would make it impossible to identify which entity resolved the vulnerability. To solve this, we used a multi-round notification design, notifying one stakeholder in each round. This allows us to identify more confidently which stakeholder resolved the vulnerability. However, there is still a risk of contamination when a stakeholder responds much later, after the next round has begun. We minimized such risks by leaving 4-5 weeks between each round of notifications, giving each stakeholder ample time to respond and remediate. Based on the email responses, we observed that nearly all arrived within two weeks of the initial notification. Additionally, publishers may reach out to ad-networks to remediate the vulnerability, before we notify ad-networks in the second phase, alerting them. However, publishers are incentivized to work with more ad-networks and vice versa to maximize their respective revenue. Hence, we acknowledge that complete isolation is not possible. However, since we detect vulnerabilities before each phase, ad-networks were notified of different or unresolved vulnerabilities than the ones in notifications to publishers, guaranteeing confidence in the results we observe. We believe these approaches effectively addressed the challenge of interdependent multi-stakeholder notification research.
@@
Source of activist notifications. We sent our activist notifications in collaboration with the CheckMyAds Institute, a well-known organization in the ad-tech community. Due to operational constraints, we could only use their branding on our emails, reports, and the website. We could not use their email servers to send notifications due to risks to their server reputation and potential interference with their operations. In theory, this could have influenced how recipients perceived the email. To minimize this impact, we clearly stated CheckMyAds’ involvement in the opening sentence of the notification and used their branding prominently in the headers and footers of our emails, reports, and website. On a positive side, results in Section 4.2 suggests email branding to play a positive role in imposing sender branding. Some publishers and advertisers acknowledged being familiar with CMA’s work, showing that email branding resulted in perception differences between activist and academic notifications—something that has not been studied before.
Other limitations. The selection of control groups based on propensity matching does not account for factors like differences in technological capabilities, resources, or internal processes between treatment and control groups. We acknowledge this to be a limitation, however, we are limited by what we can observe as outsiders to the ecosystem and cannot account for these differences. Our approach cannot identify dark pools stemming from obfuscated or encrypted advertising ID(s): also highlighted as a limitation of the approach that we follow [vekaria2023inventory].
6.3 Concluding remarks
[…]
Multi vs. single stakeholder notifications. Notifying just the highest remediating stakeholder (e.g., ad-networks in our research) via single-stakeholder notifications would have still yielded sub-optimal results. First, it is often not known apriori that which stakeholder would be the highest remediating in different multi-stakeholder ecosystems. Second, in ecosystems with interdependencies between stakeholders where incentives are mis-aligned, single-stakeholder notifications may be effective for remediation, however, influence from other stakeholders can lead to more effective remediation. Hence, notifying even just the most influential stakeholder would have proven to be sub-optimal as the most influential stakeholder (i.e., advertisers in this case) might not be the most capable stakeholder to resolve the vulnerability (i.e., ad-networks).
Implications for online advertising. Our research illuminates how the online advertising supply chain functions and how different stakeholders interact. Our qualitative analysis shows that notifications can raise awareness about vulnerabilities in the supply chain, leading to increased transparency, trust, and the termination of relationships with problematic actors. While notifications can effectively address vulnerabilities like dark pools in the online advertising ecosystem, there are important nuances. The motivations and capabilities of notification recipients play a crucial role. We found that publishers have the lowest dark pool remedial rates among stakeholders. In contrast, ad-networks, with their capability and resources, and advertisers, with their monetary incentives, are significantly more effective at resolving dark pools. Additionally, we found that the source of the notification, whether from activists or academics, does not affect dark pool resolution, except for publishers. Future work can investigate other notification channels (e.g., public naming and shaming [tingley2022effects]) and external stakeholders (e.g., direct consumer notifications [ahmad2024companies]).
Operationalizing our notification pipeline to combat industry-scale ad inventory fraud Our findings provide a clear operational strategy for industry-wide, real-time notification feasibility. IAB Tech Lab—that introduced ad-tech supply chain transparency standards—already crawls ads.txt and sellers.json files on a daily-basis for automated validation checks via their Transparency Center [iab_transparency_center]. They are in the best position to directly integrate our vulnerability detection logic based on their crawled files and by incorporating a dynamic analysis component to automatically flag and notify entities whose inventory relationships appear invalid or indicative of dark pooling. As IAB already maintains industry-wide mapping of inter-organizational relationships across both major and long-tail ad-tech participants, it can fully automate the validation and remediation loop. Furthermore, since the number of ad-networks—found to be the most influential stakeholders at mitigating dark pooling—is orders of magnitude smaller than the millions of publishers and brands, the IAB can easily maintain accurate contact channels such as verified email addresses for these entities, enabling targeted, high-impact notifications at scale. In combination, these capabilities and our research enables IAB to deploy a continuous, industry-level fraud-prevention pipeline that is more comprehensive and effective than any isolated remediation effort.
Implications for other supply chains. Our results highlight the need to consider the complexities of multiple stakeholders in other ecosystems, such as the software supply chain, which includes operators, administrators, vendors, and developers. This need is further highlighted by Zimmermann et al. [Zimmermann2019] who showed that vulnerabilities in widely-used open-source libraries impacted numerous software products. Unfortunately, the current state-of-the-art in security and privacy vulnerability notification only considers singlular stakeholders. Similarly, in fintech, Abdou et al. [Abdou2020] showed that inadequate API security could expose critical consumer financial data, impacting multiple stakeholders including consumers, banks, developers, and regulators in open banking ecosystem. Our work provides a model for testing the effectiveness of vulnerability notifications in these multi-stakeholder ecosystems. When considering multi-stakeholder notifications in other ecosystems, we recommend identifying the most influential and the most capable stakeholders, along with the incentive flows and interdependencies within the ecosystem. Our work guides recommendation to (1) notify the most capable stakeholder prior to most influential one, and (2) the most interdependent stakeholder before the less interconnected ones. This is in order to create a build-up of pressure on the entities notified in latter stages.
Multi-stakeholder notification research warrants more attention. Research on multi-stakeholder supply chains is complicated due to stakeholder inter-dependencies, existence and fixing capabilities of vulnerabilities across stakeholders and with mixed incentives. These complexities raise questions about which stakeholder to notify and how to notify each one. Our study, using the online advertising supply chain as a case-study, shows that not all stakeholders are equally capable or willing to resolve vulnerabilities. Furthermore, changing notification sources only, significantly influences some stakeholders. This research is the first to tackle this scenario and highlights the need and aids examination of other fragmented multi-stakeholder ecosystems.
7 Ethical Considerations
@@
IRB review. Our institutional review board (IRB) reviewed this notification study and deemed it as not involving human subjects. Despite this, we designed and conducted our measurements and notifications following the principle of beneficence outlined in the Menlo Report [MenloReport2012] and Belmont Report [BelmontReport1979]. We aimed to maximize benefits and minimize potential harms.
@@
Infrastructure costs and risks. We measured ad-tech supply vulnerabilities through crawls. To avoid stressing web servers, we did not conduct dynamic crawls of problematic publishers concurrently. We spaced out our additional crawls for ads.txt and sellers.json files between 12 days to one month apart. Our crawlers did not follow the robots.txt directives on problematic publishers, but our methodology aligns with ethical and legal considerations for such audits [sandvig2014auditing]. We sent notifications from a dedicated email server, set up with our university’s IT staff, to protect university mail server’s reputation. We did not send repeated probes or marketing-style messages. The content was minimal—not the actual notification content, but just one sentence—asking to schedule a zoom call and we received no positive responses, ensuring negligible inbox burden on the recipients.
Participation and Privacy risks. We neither collect or record any personal information in this research nor include any organization names in the paper. Although we had phone calls and virtual meetings with several organizations to help them remediate vulnerabilities, we did not record or share any conversations or emails with anyone other than the authors. Moreover, we respected opt-out decisions of participants and deleted all their data. Using a codebook to code responses helped us prevent any deanonymization of the source of the response. We are committed to not sharing the actual mapping of the ad-tech entity name and their responses. We described our study to be part of a research study, disclosed our affiliations, and clearly outlined potential benefits and risks from the reported vulnerabilities in our notifications as well as on the project website shared with the recipients.
@@
Advertising costs. To understand which brands advertise on problematic publishers and which ad-networks are responsible, we clicked on the ads shown during page loads. The costs associated with our 28,376 ad clicks across 12 crawls of problematic publishers represented by label C and C’ in Figure 3 of the paper were relatively minor, as per-ad CPMs are low, particularly for browsers without profiles [olejnik2013selling, cook2020inferring]. These costs are justifiable given the benefits of understanding and remedying supply-chain vulnerabilities in the ad-tech ecosystem. Quantitatively, 28,376/12 = 2365 ad clicks were performed per crawl distributed across 684 problematic publishers — 2365/684 $\sim$3.5 ad clicks per publisher on an average. The average CPC costs for the open programmatic ad-tech ecosystem is \$0.63 for 2024 [google_cpc_rates_2024] (as estimated for the dominant player – Google). So, the maximum revenue benefit our study could have caused using the average case is 3.5 * 0.63 = \$2.2 per problematic publisher.
@@
Delayed notification to Control entities We adopted delayed notification for control entities by notifying them at the end of the study only because it was the minimal-risk design that allowed sound attribution of the remediation effects. No alternative could have produced valid results without substantially increasing methodological (Section 3.2) and ethical risks. Notifying controls earlier—either at the time vulnerabilities were first computed or immediately after each round—would have collapsed the distinction between treatment and control. This is because notifying control entities could have resulted in them reaching out to other ad-tech entities for resolution, making it difficult to understand if the remediation observed in the current as well as the following round was due to treatment notification or due to potential poisoning of some highly interconnected control entities (via business relationships) with treatment entities. At the same time, the actual harms from this delay were objectively minimal. For publishers and ad-networks, dark pooling creates no monetary loss, and reputational risk is inherently constrained because adversaries disperse misuse across many IDs to avoid detection; for advertisers—the highest-risk group—we explicitly placed them in the shortest delay window (i.e. the last notification round), thereby minimizing exposure. From an external point-of-view, quantifying precise harm is infeasible without acces to the internal logs. However, all available evidence—including the very small number of vulnerable IDs per entity and the absence of any negative feedback from notified controls—indicate that any temporary risk was limited. Critically, our design adheres to Belmont [BelmontReport1979] and Menlo [MenloReport2012] principles: (1) it maximizes beneficence by enabling the first rigorous, multi-stakeholder evaluation of vulnerability notification efficacy in order to build automated remediation pipelines that will protect orders of magnitude more entities. (2) it distributes burdens fairly across different stakeholder types. Delayed controls are standard in randomized trials across disciplines—often with far higher stakes—and are accepted because they are the only way to produce generalizable knowledge. Consistent with those norms, we additionally mitigated residual risks by providing control entities with detailed guidance, responsive support, and post-study assistance via multiple email correspondance and/or virtual calls to ensure that they can fix the issues or clarify doubts with us throughout the process. Thus, delayed notification was the ethically preferable and methodologically necessary choice, posing only minor, temporary risks while enabling long-term, large-scale harm reduction that no other study design could have achieved.
Acknowledgments
This work was supported in part by the National Science Foundation under award numbers 2138139, 2103439, and 2338377.
Appendix A Data Availability
We will open source our dark pooling vulnerability detection and notification pipeline upon acceptance of this work. We will also include our notification template, vulnerability remediation guidance shared with the vulnerable entities and the website we used for the notified entities to access this information. We hope this will foster future notification studies as well industry initiatives such as IAB to incorporate such notification pipelines at larger scale.
Appendix B Example: ads.txt, sellers.json, Dark Pooling
Figure 5 represents structure of sample entries in ads.txt and sellers.json files. It also illustrates a typical dark pooling scenario.
Figure 5: An example depicting the structure of entries in hypothetical ads.txt and seller.json files. In this example, publishers A, B, and C are considered unrelated (i.e., are not organizational related) and represents a scenario of dark pooling. The sellers.json of adxA.com shows that the seller account 12345 is owned by PublisherC. However, ads.txt of A.com and B.com lists seller account 12345 on adxA.com as their direct seller. This suggests dark pooling where A and B are pooling their ad inventories with C on adxA.com.
Appendix C Categorization of Notification Responses
@@
Table V shows categorization of notification responses by sender reputation under different themes and thematic categories.
TABLE V: Categorization of notification responses by sender reputation under different themes and thematic categories for victim publishers, ad-networks, and advertisers.
Theme Thematic Catgeories Victim Publishers Ad Networks Advertisers
Activist Academic Activist Academic Activist Academic
Approaches to Resolution Actionable details requested 9 5 6 7 0 1
Forwarded to the concerned team 7 10 0 0 5 0
Collaborative resolution performed 3 2 4 4 0 1
Fixed the reported issues 7 2 1 0 1 0
Contacted responsible entities for fix 3 3 0 2 0 1
Justifying the reported issue 1 0 2 0 0 0
Attitude towards Notification Acknowledged the notification 16 13 5 8 3 2
Motivated to fix the issue 9 5 2 3 0 2
Concerned about the notification 4 3 2 1 2 3
Research initiative appreciated 4 1 3 1 0 0
Awareness about the vulnerability 3 0 0 1 0 0
Notification unhelpful or uninterested 1 2 0 1 0 0
Challenges to Resolution Doubting correctness of the report 5 2 4 2 0 0
Misinterpreted as phishing 2 1 1 3 1 1
Limited resources to fully resolve 1 2 0 1 0 0
Complete resolution beyond control 0 1 1 1 0 0
Appendix D Propensity-matched DiD Algorithm
Algorithm 1 describes the pseudocode for performing propensity-matched difference-in-differences (PM-DiD).
Algorithm 1: Propensity Matched Difference-in-Differences
1:Input:
2:   Treatment group $\mathcal{T}$ (i.e., $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ or $\mathcal{T}_{1}\cup\mathcal{T}_{2}$)
3:   Control group $\mathcal{C}$
4:   Measure $M$
5:Output:
6:   Effect size $\Delta_{(t,c)}$
7:   Statistical measures $N_{\text{rem.}},\mu_{\text{ov.}},\mu_{\text{rem.}}$
8:Rounds and Measures:
9:   Round 1: $M\leftarrow probdomains_{vp}$
10:   Round 2: $M\leftarrow pools_{ad}$ and $partnerpools_{ad}$
11:   Round 3: $M\leftarrow pools_{ad}$
12:
13:Step 1: Propensity Matching for each round
14:for all $t\in\mathcal{T}$ do
15:  Compute measure: $M^{t}_{\text{before}}$
16:  Find nearest neighbor $c\in\mathcal{C}$ such that:
17:    $\min(|M^{t}_{\text{before}}-M^{c}_{\text{before}}|)$
18:  Add matched pair $(t,c)$ to matched pairs $\mathcal{P}$.
19:end for
20:
21:Step 2: Measure evaluation before and after intervention
22:for all $(t,c)\ pairs\in\mathcal{P}$ do
23:  Record pre-intervention measure values: $M^{t}_{\text{pre}}$, $M^{c}_{\text{pre}}$
24:  Record post-intervention measure values: $M^{t}_{\text{post}}$, $M^{c}_{\text{post}}$
25:  Compute effect size per pair:
26:    $\Delta_{(t,c)}=(M^{t}_{\text{post}}-M^{c}_{\text{post}})-(M^{t}_{\text{pre}}-M^{c}_{\text{pre}})$
27:end for
28:
29:Step 3: Statistical Metrics
30:Initialize counters: $N_{\text{rem.}}\leftarrow 0$, $\mu_{\text{ov.}}\leftarrow 0$, $\mu_{\text{rem.}}\leftarrow 0$
31:for all $(t,c)$ pairs do
32:  if $\Delta_{(t,c)}<0$ then
33:   Increment $N_{\text{rem.}}\leftarrow N_{\text{rem.}}+1$
34:   Update $\mu_{\text{rem.}}\leftarrow\mu_{\text{rem.}}+\Delta_{(t,c)}$
35:  end if
36:  Update $\mu_{\text{ov.}}\leftarrow\mu_{\text{ov.}}+\Delta_{(t,c)}$
37:end for
38:Compute final metrics:
39:    $\mu_{\text{ov.}}=\frac{\mu_{\text{ov.}}}{|\mathcal{T}|}$
40:    $\mu_{\text{rem.}}=\frac{\mu_{\text{rem.}}}{N_{\text{rem.}}}$
41:
42:Return: $\Delta_{(t,c)},N_{\text{rem.}},\mu_{\text{ov.}},\mu_{\text{rem.}}$